OpenStudy (anonymous):

The total value of a collection of quarters and nickels is $3.75. If the total number of coins is 27, how many of each type of coin are there?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

hero (hero):

let: n = nickels q = quarters, then: n + q = 27 .05n + .25q = 3.75 Solve the system

OpenStudy (anonymous):

75 nickels

OpenStudy (anonymous):

@Hero

hero (hero):

There's only 27 coins in total. So there can't be 75 nickels. Do you know how to solve systems of equations?

OpenStudy (anonymous):

nope

hero (hero):

This video should help: https://www.youtube.com/watch?v=nok99JOhcjo

OpenStudy (anonymous):

3.75 - .25=3.50

OpenStudy (imstuck):

Are you still having trouble with this?

OpenStudy (anonymous):

yea

hero (hero):

@alyygirl, the goal is to solve for variables n and d.

OpenStudy (imstuck):

Ok, first relate the NUMBER of coins in one equation, and then in the second equation, deal with the VALUE of the coins. You have nickels, n, and quarters, q, and there is a total number of 27 of them, so the first equation is: q + n = 27.

OpenStudy (anonymous):

well the answer would help since solving it makes no sense

OpenStudy (imstuck):

The second equation will relate the value of the coins to the total value you're given which is 3.75. The difficult thing to do here is to remember that since a quarter is worth .25, then the value of the number of quarters you have, q, is .25q. Do the same with the nickels. The number of nickels you have, n, translates to .05n, because each nickel you have, even though you don't know the number of nickels, is worth 5 cents. This value of quarters, .25q, addded to the value of nickels, .05n equals 3.75. So the second equation is .25q + .05n = 3.75. Both equations look like this now:

OpenStudy (imstuck):

q + n = 27 .25q + .05n = 3.75

OpenStudy (imstuck):

Do you know how to solve a system of equations?

OpenStudy (anonymous):

im stuck from there actually

OpenStudy (imstuck):

Hint: solve the first equation for q. Like this... q =

OpenStudy (anonymous):

how would I get q?

hero (hero):

Start with the equation q + n = 27 Then subtract n from both sides.

OpenStudy (imstuck):

Subtract n from both sides, like this: q + n - n = 27 - n On the left side, a +n and a -n cancel each other out. So what does q= ?

OpenStudy (imstuck):

Youre not solving for a value for q when we ask you to solve that equation for q, we mean to put the equation in terms of q. That means, "q = " That's in terms of q.

OpenStudy (anonymous):

there is no a

OpenStudy (imstuck):

No a, that was meant as a "+n" and a "-n". +n -n = 0. N's are gone.

OpenStudy (anonymous):

23.25

OpenStudy (imstuck):